### v1

# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

Abstract
Causal inference typically assumes centralized access to individual-level data. Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints. We address this by estimating the Average Treatment Effect (ATE) from decentralized observational data using federated learning, which enables inference through the exchange of aggregate statistics rather than individual-level data. We propose a novel method to estimate propensity scores in a (non-)parametric manner by computing a federated weighted average of local scores, using two theoretically grounded weighting schemes—Membership Weights (MW) and Density Ratio Weights (DW)—that balance communication efficiency and model flexibility. These federated scores are then used to construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW). Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap. We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions, with theoretical analysis and experiments on simulated and real-world data highlighting their strengths and limitations relative to meta-analysis and related methods.
1 Introduction
In Randomized Clinical Trials (RCTs), treatment assignment is randomized, which ensures that the observed association between treatment and outcome reflects a causal effect. Under this design, the Average Treatment Effect (ATE) can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman et al.,, 1990), which can be further refined through covariate adjustment to reduce variance (U.S. Food and Drug Administration,, 2023; European Medicines Agency,, 2024; Lei and Ding,, 2021). However, RCTs are often expensive, time-consuming, or infeasible in practice. In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán,, 2018; Hernán and Robins,, 2006). Although real-world data is abundant, drawing causal inferences from it remains challenging due to the presence of confounding covariates. As a result, the unadjusted DM estimator is biased when estimating the ATE (Grimes and Schulz,, 2002). To mitigate this bias, it is essential to adjust for confounding variables (VanderWeele,, 2019). This can be done by predicting the counterfactual outcomes of the individuals under the opposite treatment before averaging the differences between the observed and counterfactual outcomes, leading to the so-called G-Formula plug-in estimator (Robins,, 1986). Another popular method for deconfounding is to weight individuals according to their probability of receiving the treatment, thereby emulating a randomized experiment. The Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin,, 1983) relies on the estimation of the propensity score—the probability of treatment given the covariates. Doubly robust estimators, such as the Augmented IPW (AIPW) (Bang and Robins,, 2005), enhance IPW by incorporating an outcome model, offering robustness against misspecification of either model.
In causal inference, larger datasets enhance the precision and reliability of treatment effect estimates, particularly for underrepresented subgroups often overlooked in smaller studies. However, real-world data is typically decentralized—collected across several hospitals, companies, or countries—making it challenging to aggregate into a centralized dataset. This difficulty is particularly acute in domains like healthcare, where privacy concerns, regulatory barriers, data ownership and logistical issues (such as responsibility for data storage and governance) complicate data sharing. Federated Learning (FL) (Kairouz et al.,, 2021) offers a promising solution by enabling model training directly on decentralized data within a server-client architecture, without sharing individual-level data. By exchanging only aggregate statistics such as model updates or gradients, FL has the potential to support collaborative causal inference studies with performance approaching that of centralized analyses. However, realizing this potential requires tailored methodological developments.
Contributions. In this article, we build on ideas from Federated Learning to design (A)IPW estimators for decentralized observational data, moving beyond the simple aggregation of local ATE estimates used in traditional meta-analysis approaches.
At the core of our approach is a flexible, nonparametric strategy for federating propensity scores. Unlike prior methods that aggregate parameters of a global parametric model via “one-shot” averaging (Xiong et al.,, 2023) or federated gradient descent (Khellaf et al.,, 2025), we construct a global propensity score model as a mixture of locally estimated models. This is done through a two-step process: (1) local estimation of propensity scores at each site, which naturally accommodates site-specific heterogeneity in treatment assignment mechanisms and offers flexibility in model choice, and (2) aggregation of these local models into a global one using carefully selected federated weights. We introduce two theoretically grounded weighting schemes: Membership Weights (MW), which represent the probability of an individual belonging to each site given their covariates, and Density Ratio Weights (DW), which model the relative density of covariates at each site compared to the global population. We explain how to infer MW and DW weights in a federated manner and highlight the distinct trade-offs they offer between communication efficiency and modeling assumptions. Using these federated propensity scores, we construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW). We derive asymptotic variance expressions for these estimators and compare them to the meta-analysis estimators—often overlooked in FL (Khellaf et al.,, 2025)––, proving that our estimators achieve lower or equal variances in all scenarios considered.
Our approach is particularly beneficial in scenarios with poor or nonexistent overlap between treated and control groups within individual sites. In such cases, cross-site collaboration becomes crucial, as propensity score functions vary across sites. Indeed, when treatment assignment mechanisms differ substantially—such as when one site treats a subgroup that remains untreated elsewhere—the global dataset achieves markedly greater overlap, providing essential support for estimating treatment effects that would otherwise be poorly estimated within isolated sites. Additionally, our framework naturally accommodates a wide range of cross-site heterogeneities, including variations in sample sizes, treatment policies, covariate distributions, and violations of positivity. Numerical experiments on both simulated and real-world data validate our theoretical findings and highlight the practical value of our method.
2 Preliminaries
2.1 ATE estimators from centralized multi-site observational data
In this section, we recall the key components of ATE estimation in a centralized multi-site setting. Following the potential outcomes framework (Rubin,, 1974; Splawa-Neyman et al.,, 1990), we consider random variables $(X,H,W,Y{(1)},Y{(0)})$, where $X\in\mathbb{R}^{d}$ represents patient covariates, $H\in[K]$ indicates site membership, $W\in\{0,1\}$ denotes the binary treatment, and $Y(1)$ and $Y(0)$ are the potential outcomes under treatment and control, respectively. We assume that the Stable Unit Treatment Values Assumption (SUTVA) holds, so that the observed outcome is $Y=WY(1)+(1-W)Y(0)$, and that potential outcomes are uniformly bounded.
In the centralized setting, we observe $n=\sum_{k=1}^{K}n_{k}$ observations of independently and identically distributed (i.i.d.) tuples $(H_{i},X_{i},W_{i},Y_{i})_{i\in 1,\dots,n}\sim\mathcal{P}$, with $n_{k}=\sum_{i=1}^{n}\mathds{1}_{\{H_{i}=k\}}$ the number of observations in site $k$. We aim to estimate the ATE defined as the risk difference $\tau=\mathbb{E}\left[Y(1)-Y(0)\right]=\mathbb{E}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$, where the expectation is taken over the population $\mathcal{P}$. To be able to identify the ATE, we assume unconfoundedness.
Assumption 1 (Unconfoundedness).
$(Y(0),Y(1))\perp\!\!\!\perp W\mid X$.
While Assumption 1 is a standard assumption in causal inference, we further consider 2, which is specific to the multi-site setting. Together with Assumption 1, it ensures that the variables $X$ form a sufficient set of covariates to adjust for confounding.
Assumption 2 (Ignorability on sites).
$(Y(0),Y(1))\perp\!\!\!\perp H\mid X$.
Our setting is depicted in the graphical model in Figure 1, which highlights that we eliminate any direct effect of the site on the outcome.
$W$$X$$Y$$H$
Figure 1: Graphical model for multi-site observational data.
We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$. The oracle propensity score is denoted by $e(x)=\mathbb{P}(W\mid X=x)$, and we consider the weak (global) overlap assumption (Wager,, 2024).
Assumption 3 (Global overlap).
$\mathcal{O}_{\mathrm{global}}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\big{]}<\infty$.
Assumption 3 is crucial for propensity score-based estimators, as it states that every region of the covariate space has a non-zero probability of receiving both treatments. A lower value of $\mathcal{O}_{\mathrm{global}}$ indicates that these probabilities lie further away from 0 and 1, which corresponds to better overlap. For further insights on overlap, see Li et al., 2018a ; Li et al., 2018b , and for a “misoverlap” metric, refer to Clivio et al., (2024).
With Assumptions 1, 2 and 3, the ATE is identifiable as $\tau=\mathbb{E}\big{[}\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\big{]}$ (see Appendix A.1). We denote oracle ATE estimators, which assume knowledge of the nuisance components $e,\mu_{0},\mu_{1}$, by a superscript . We define the centralized, multi-site estimators as follows.
Definition 1 (Oracle multi-site centralized estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),\quad\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (1)
where $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$ and $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}$.
These oracle estimators are unbiased and asymptotically normal.
Theorem 1 (Consistency of oracle multi-site centralized estimators).
Under Assumptions 1, 2 and 3, we have the following asymptotic normality results:
$\sqrt{n}(\hat{\tau}-\tau)\to\mathcal{N}(0,V^{\star})\ \text{with}$
$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{*}\right]=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},$
$\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{*}\right]=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right],$
where $\tau(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ denotes the Conditional Average Treatment Effect (CATE).
These asymptotic variances align with those in the single-site setting (Hirano et al.,, 2003), as detailed in Appendix A.2. However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data. This creates a challenge in the decentralized setting, where centralizing data to compute $\mu_{1}$, $\mu_{0}$, and $e$ is not feasible. Therefore, the estimators in Definition 1 need to be adapted to this setting. Importantly, the AIPW estimator is doubly robust, meaning it remains consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al.,, 2018).
2.2 Meta-analysis estimators
We now turn to a decentralized setting in which the $K$ sites cannot share individual-level data. A natural baseline for estimating the ATE across sites is a two-stage meta-analysis approach (Burke et al.,, 2017), wherein each site independently estimates the relevant nuisance parameters and communicates only the resulting ATE estimates for aggregation. We will need the following assumption.
Assumption 4 (Local overlap).
$\forall k\in[K],\mathcal{O}_{k}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\mid H=k\big{]}<\infty$.
Assumption 4 is much stronger than its global counterpart (Assumption 3), as it requires the overlap to hold locally for each site. Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows.
Definition 2 (Oracle meta-analysis estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (2)
where $\hat{\tau}_{\mathrm{IPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e_{k})$ and $\hat{\tau}_{\mathrm{AIPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e_{k},\mu_{1},\mu_{0})$ are the local estimators at site $k$.
While alternative aggregation weights—such as the inverse variance of local estimates—can be considered, they produce, in our setting, biased estimates of the global ATE $\tau$, which decomposes as $\tau=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, where $\rho_{k}=\mathbb{P}(H=k)$ and $\tau_{k}=\mathbb{E}\left[Y(1)-Y(0)\mid H=k\right]$ denotes the local ATE at site $k$. This bias arises whenever the $\{\tau_{k}\}_{k}$ differ, which commonly occurs when covariate distributions vary across sites and treatment effects are heterogeneous (i.e., depend on covariates).
Theorem 2 (Consistency of oracle meta-analysis estimators).
Under Assumptions 1, 2 and 4, the oracle meta-analysis estimators are unbiased for the ATE with asymptotic variances
$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right],\quad\quad\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right]$
with within-site variances $V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}$, $V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right]$ and between-sites variance of the local ATEs $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$.
This result is proved in Appendix A.3. A key limitation of meta-analysis estimators is their reliance on Assumption 4, which is fragile and often violated in practice—for example, when a site follows a deterministic treatment policy (e.g., treating all patients or only those with a specific condition). In such cases, these estimators become ill-defined and yield biased estimates of the global ATE.
To overcome this, we introduce a federated approach that estimates the global propensity score $e$ as a weighted combination of the local scores $e_{k}$, enabling valid inference even without local overlap.
3 Related work
Federated causal inference has only recently gained attention. Khellaf et al., (2025) propose a federated framework for estimating the ATE across multiple RCTs by fitting separate parametric linear outcome models for treated and control groups at each site using federated learning, followed by ATE estimation via the G-formula. Their method is motivated by the fact that, in RCTs, this estimator remains unbiased under model misspecification and achieves lower variance than the Difference-in-Means estimator.
In observational studies, however, outcome model misspecification can bias ATE estimates. To address this, Vo et al., (2022) introduce a Bayesian framework using Gaussian processes trained via federated gradient descent. Their method requires estimating a multi-site covariance kernel, which involves each site computing and sharing the first four moments of its data—a computationally intensive process that raises concerns regarding scalability and communication efficiency.
Alternative approaches, such as IPW, require estimating the propensity score. Guo et al., (2025) propose a federated method for learning a global propensity score model using a consensus voting scheme over parametric model parameters, retaining only sites that conform to a shared specification. In contrast, our approach does not assume a common propensity score function: each site can adopt its own, potentially distinct model.
To address site-level heterogeneity, Xiong et al., (2023) assume a parametric logistic propensity score model with shared and site-specific parameters, federating only the common ones in a single round of communication round. In a similar line of work, Yin et al., (2025) estimate a global parametric propensity score model that adjusts for both covariates and site membership, but their handling of heterogeneity is limited to a site-specific scalar, restricting the model’s flexibility. In contrast, our method makes no structural assumptions about the form of the propensity score. We enable fully nonparametric global estimation while allowing flexible and heterogeneous local models, without requiring prior knowledge of shared or site-specific components. Crucially, unlike the above methods, which assume overlap holds locally at each site, our approach explicitly relaxes this strong assumption.
Although not always framed within a federated context, a related body of work focuses on the generalization of causal findings from source data to a target population. Assuming homogeneous outcome models and propensity scores across sites, Han et al., (2025) propose a density ratio weighting method to generalize ATE estimates from multiple source sites to a target domain. Their approach resembles meta-analysis, relying on aggregate statistics—such as sample sizes and local ATE variances—rather than individual-level data. While they estimate density ratios using parametric models, our Membership Weights approach leverages supervised learning methods (e.g., logistic regression) that have lower sample complexity and make less stringent assumptions on the sites’ distributions. Similarly, Guo et al., (2024) address generalization to a target population composed of multiple source sites and previously excluded individuals. Their method assumes the target propensity score is proportional to those of the sources and estimates local scores via density ratios between treated or control subgroups and the target population. However, nonparametric estimation of these ratios requires sufficient data in each treatment arm at every site—a condition often unmet in small or imbalanced settings. Furthermore, the authors do not address how this could be implemented in a federated setting.
4 Federated estimators via propensity score aggregation
4.1 Oracle federated estimators
As discussed in the previous section, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al.,, 2025), site differences limited to intercept shifts (Yin et al.,, 2025), or predefined shared structures (Xiong et al.,, 2023). In practice, treatment assignment often varies across sites due to differences in norms, resources, or clinical practices. To accommodate this heterogeneity, we express the global propensity score as a weighted combination of site-specific scores, capturing local treatment policies.
Specifically, the global propensity score can be expressed as a weighted combination of local scores using two alternative weighting schemes (see Appendix A.4). The first relies on Membership Weights (MW), representing the probability of belonging to each site given the covariates:
$e(X)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X)}_{\text{membership weights}}e_{k}(X).$ (3)
The second uses Density ratio Weights (DW), based on the ratio between the covariate density at site $k$, denoted $f_{k}$, and the overall population density $f$:
$e(X)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(X)}{f(X)}}_{\text{density weights}}e_{k}(X).$ (4)
Building on Eq. 3 and 4, we define our oracle Federated (A)IPW estimators, denoted Fed-(A)IPW, where the corresponding weights will be estimated via federated learning (see Section 4.2).
Definition 3 (Oracle federated estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}},$ (5)
where $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e)$, and $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})$ rely on the global propensity score $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$ with federated weights $\{\omega_{k}(X)\}_{k}$ that are either the membership (Eq. 3) or the density ratio (Eq. 4) weights.
Crucially, our decompositions of the global propensity score in Eq. 3 and 4 enable a flexible combination of locally estimated propensity scores $e_{k}$ and globally learned federation weights $\omega_{k}(x)$. Theorem 3 (proved in Appendix A.5) establishes that this approach, in the oracle setting, matches the efficiency of centralized estimators.
Theorem 3 (Equality of centralized and federated (A)IPW estimators).
Under Assumptions 1, 2 and 3, the oracle federated estimators (Def. 3) are equal to the oracle centralized estimators (Def. 1).
We further show in Appendix A.6 that even when local overlap (Assumption 4) holds, our federated estimators achieve lower variance than meta-analysis estimators.
Theorem 4 (Variance comparison between meta-analysis and federated estimators).
Under Assumptions 1, 2 and 4, both federated and meta-analysis estimators can be computed, and
$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right],$
$\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right],$
with equality when the local propensity scores are equal.
We provide two main reasons for this improvement. First, our decomposition of $e$ in Eq.3 marginalizes over $H\mid X$, effectively removing the need to adjust for site membership $H$. This reduces variance by avoiding unnecessary conditioning. Second, our federated approach inherently yields better overlap than the siloed meta-analysis framework, as formalized in the following theorem.
Theorem 5 (Overlap improvement).
$0\leq\mathcal{O}_{\mathrm{global}}\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}.$
Theorem 5 (proved in Appendix A.7) establishes that global overlap is always at least as good as the worst local overlap. This implies that even when local overlap holds, sites with poor overlap still benefit from the federated approach: the global propensity scores $e(X)$ are more bounded away from 0 and 1 than the local scores $\{e_{k}(X)\}_{k\in[K]}$. Notably, sites with poor individual overlap can even improve the overall overlap of the federated population, as illustrated in the following example.
Example 1. Let $K=2$ with $X_{i}=1$ in both sites, $n_{1}=n_{2}$, $\mathbb{P}(H_{i}\mid X_{i})=0.5$, $e_{1}(X_{i})=0.99\times X_{i}$ and $e_{2}(X_{i})=0.01\times X_{i}$, leading to $e(X_{i})=0.5\times e_{1}(X_{i})+0.5\times e_{2}(X_{i})=0.5\times X_{i}$. Then, $\mathcal{O}_{1}=\mathcal{O}_{2}=\left(0.99\times 0.01\right)^{-1}\approx 101$, while $\mathcal{O}_{\mathrm{global}}=\left(0.5\times 0.5\right)^{-1}=4$. This illustrates how heterogeneity among local propensity scores over a shared population can enhance global overlap.
4.2 Federated estimation
We now move beyond oracle estimators and describe how to implement the Fed-(A)IPW estimators in a practical federated learning setting. Our method follows a two-step procedure (which can be executed in parallel): each site $k$ first estimates and shares a local propensity score function $\hat{e}_{k}(x)$; then, global scores $\hat{e}(x)$ are computed as weighted averages of the local scores, using weights $\{\hat{\omega}_{k}(x)\}_{k\in[K]}$ estimated in a federated manner. For Fed-AIPW, a third step is needed to train the outcome models $\hat{\mu}_{0},\hat{\mu}_{1}$ with federated learning. Details on how to estimate $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$, along with $\hat{\mu}_{0},\hat{\mu}_{1}$, are provided in the following subsections.
4.2.1 Estimation of the local propensity scores
Local propensity scores can be estimated using any probabilistic binary classification method. Common choices include logistic regression for parametric estimation and generalized random forests for non-parametric estimation (Lee et al.,, 2010). A key advantage of our approach is its flexibility: different estimation methods can be used across sites, tailored to local data characteristics or computational constraints.
As previously discussed, this procedure does not require the overlap assumption to hold locally. That is, local propensity scores can legitimately approach 0 or 1—as long as the global score remains sufficiently bounded away from these extremes. This property makes the approach suitable for scenarios involving external control arms (Center for Drug Evaluation and Research and Center for Biologics Evaluation and Research and Oncology Center of Excellence,, 2023; European Medicines Agency,, 2023) where some sites have $e_{k}(X_{i})=0$ for all $i$ in the control group, yet still contribute to the global analysis.
4.2.2 Estimation of the federated weights
Membership weights.
We begin with membership weights $\{\omega_{k}(x)=\mathbb{P}(H_{i}=k\mid X=x)\}_{k}$, which represent the probability of site membership given covariates. These weights can be estimated without access to individual-level data using any federated algorithm for multi-class classification models that can output probabilities.
Federated versions of parametric models—such as logistic regression or deep neural networks—trained via gradient-based empirical risk minimization are well-suited for this task and readily supported in practice (Kairouz et al.,, 2021). Additionally, non-parametric models like random forests (Hauschild et al.,, 2022) and gradient-boosted trees (Li et al.,, 2020) have been adapted to the federated setting, offering more flexibility in capturing complex decision boundaries.
In our experiments, we use a simple multinomial logistic regression model trained via the standard Federated Averaging (FedAvg) algorithm (McMahan et al.,, 2017). This approach is easy to deploy in a server–client architecture, remains interpretable, and requires exchanging only $TKd$ floats, where $T$ is the number of communication rounds. We refer to Appendix B for details.
Density ratios weights.
We now turn to density ratio weights $\{\omega_{k}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}\}_{k}$, where $f_{k}$ and $f$ denote the local and global covariate densities, respectively. Although various methods exist for estimating density ratios (Sugiyama et al.,, 2012), we are not aware of any that are directly compatible with federated settings. Standard non-parametric techniques, such as kernel density estimation, typically require sharing raw data, which violates the decentralized constraints we impose.
We thus propose to estimate each local density $f_{k}$ using a parametric model (e.g. via maximum likelihood estimation). In our experiments, we assume that covariates are normally distributed within each site. Each site $k$ locally estimates its mean $\hat{\mu}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}X_{i}$ and covariance matrix $\hat{\Sigma}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}(X_{i}-\hat{\mu}_{k})(X_{i}-\hat{\mu}_{k})^{\top}$, then transmits these to the central server. Noting that the global density $f$ is a mixture of the $K$ local densities weighted by site proportions $\rho_{k}$ (which can be estimated by $n_{k}/n$), we estimate the density ratios weights as $\hat{\omega}_{k}(x)=\frac{\frac{n_{k}}{n}\hat{f}_{k}(x)}{\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{f}_{k}(x)}$, where each local density is given by
$\hat{f}_{k}(x)$ $=\frac{1}{(2\pi)^{d/2}|\hat{\Sigma}_{k}|^{1/2}}\times\exp\Big{(}-\frac{1}{2}(x-\hat{\mu}_{k})^{\top}\hat{\Sigma}_{k}^{-1}(x-\hat{\mu}_{k})\Big{)}.$
In contrast to the estimation of membership weights, this approach requires only a single round of communication, but involves sending $Kd+Kd^{2}$ floats, which can be significant for high-dimensional data.
More broadly, we recommend using a parametric approach to density ratio estimation when prior knowledge about the local distributions is available and well-justified. However, in the absence of reliable distributional assumptions, membership weights often offer a more robust and practical alternative. They allow for flexible parameterizations and can be estimated through classification models, which are generally easier to learn than densities—particularly in high-dimensional settings where density estimation suffers from high sample complexity and poor scalability (Sugiyama et al.,, 2012).
4.2.3 Estimation of Fed-AIPW
Fed-AIPW requires estimating the outcome models $\mu_{1}$ and $\mu_{0}$ in addition to the propensity components $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$. The estimation procedure for $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$ follows the same steps outlined in Sections 4.2.1 and 4.2.2. For the outcome models, one can estimate $\mu_{0}$ and $\mu_{1}$ globally using federated gradient descent, as in Khellaf et al., (2025), with parametric models such as logistic regression for binary outcomes or linear regression for continuous outcomes. As with the estimation of membership weights, our framework also supports more flexible outcome modeling using a broad range of function classes, including neural networks, decision forests, and other gradient-based models adapted to the federated setting.
5 Experiments
Synthetic data.
We consider $K=3$ sites and $d=10$ covariates. Two distinct data-generating processes (DGPs) are used in the simulations. In DGP A, each site $k$ independently samples $n_{k}=500$ individuals from a site-specific multivariate Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$. In DGP B, a total of $n=1500$ individuals are first drawn from a bimodal Gaussian mixture and then assigned to sites using a multinomial logistic model based on their covariates. We vary the degree of overlap within sites to reflect different practical scenarios: No local overlap (the second site has no treated individuals, hence $\mathcal{O}_{2}=\infty$), Poor local overlap ($\mathcal{O}_{2}\approx 10^{7}$), and Good local overlap ($\mathcal{O}_{2}\approx 4.6$). The outcome models $\mu_{1}$ and $\mu_{0}$ are shared across sites and specified as linear functions with interaction terms. For comparison, we also generate data consistent with the setting described in Xiong et al., (2023) (Figure 5). All reported results are aggregated over 1,500 simulation runs. Complete details of the simulation setup are provided in Appendix C.
We evaluate several estimators: (i) our proposed Fed-IPW and Fed-AIPW, with either Membership Weights (MW) (Def. 3) estimated via federated multinomial logistic regression (well-specified in DGP B), or Density Ratio Weights (DW) (Def. 4) based on Gaussian models (well-specified in DGP A); (ii) the Centralized Oracle (Def. 1); (iii) meta-analysis IPW and AIPW estimators with sample size weighting (Meta-SW) (Def. 2); and (iv) the one-shot inverse variance weighted IPW estimator (1S-IVW) from Xiong et al., (2023), evaluated under a favorable setting for this method—specifically, we assume that certain propensity score parameters are shared across sites. For all estimators, propensity scores are estimated with logistic regression, and outcome models with linear regression.
The results for the three degrees of local overlap are shown in Figures 5, 5 and 5, with the top (respectively bottom) row showing DGP A (respectively DGP B). Before discussing each setting separately, we start by some general observations. First, Fed-IPW-DW is unbiased in DGP A but biased in DGP B, while our Fed-IPW-MW is unbiased in DGP B but biased in DGP A, as expected. Note that Fed-IPW-MW could be made unbiased by using more flexible models (e.g., neural networks) to estimate membership probabilities. Second, our Fed-AIPW estimators remain unbiased across all overlap levels thanks to their double robustness property: despite misspecified linear outcome models, well-estimated propensity scores ensure consistency. Finally, in most cases our Fed-IPW estimators exhibit lower variance than the oracle IPW estimator. This is because they leverage estimated global propensity scores, a strategy known to reduce variance in IPW estimation (Hirano et al.,, 2003).
In the No local overlap setting (Figure 5), meta-analysis estimators are undefined due to the absence of treated individuals at one site. In contrast, our federated estimators remain unbiased under both DGPs, as Assumption 3 holds (with global overlap $\mathcal{O}_{\mathrm{global}}\approx 6.22$). In the Poor local overlap setting (Figure 5), site 2 lacks sufficient overlap, leading to bias and instability in the meta-analysis estimators, including Meta-AIPW as both the propensity weights and local outcome models are inaccurate. However, this issue is mitigated in the global dataset (see Figure 5 in the Appendix), allowing our federated estimators to perform reliably. Finally, in the Good local overlap setting (Figure 5), while all competitors are unbiased, our federated methods exhibit lower variance.
In Figure 5, we consider a specific setting in which all local propensity score functions share a common subset of 5 out of 10 logistic regression coefficients with data generated from DGP B. This setup aligns with the assumptions of the 1S-IVW method proposed by Xiong et al., (2023), which relies on prior knowledge of the shared coefficients—an assumption our method does not require. Nevertheless, Fed-MW is unbiased and exhibits the smallest variance.
Figure 2: Synthetic data: DGP A (top panel) and B (bottom).
(a) No local overlap
(b) Poor local overlap
(c) Good local overlap
Figure 3: Comparison to Xiong et al., (2023)
Figure 4: Results of Traumabase experiment.
Real data.
We apply our methods to the multi-site observational Traumabase cohort, which contains data on patients with traumatic brain injury (Mayer et al.,, 2020; Colnet et al.,, 2024). We study the effect of tranexamic acid (TA), a drug used to reduce bleeding, on the binary outcome of mortality among trauma patients. For this analysis, we select the $K=4$ centres with highest number of treated patients (min. 47), totaling 472 treated and 5,531 control patients, each with 17 measured covariates. We standardize the covariates to have zero mean and unit variance in a federated manner by computing and sharing the means and variances of the covariates across sites.
We focus on AIPW estimators. For Fed-MW and Fed-DW, we estimate the $e_{k}$’s with local logistic regression, while the outcome models $\mu_{1},\mu_{0}$ and the membership weights are learned with a logistic regression model trained with FedAvg (Algorithm 1 in the Appendix) with $T=5000$ rounds, $E=1$ local steps and learning rate $\eta=0.1$. The density ratio weights are computed under a Gaussian model for the covariates. We compare our estimators to Meta-SW and centralized AIPW estimators using also linear and logistic models. To construct empirical confidence intervals, we perform 1000 bootstrap resamples. The results in Figure 4 demonstrate that our federated approach yields estimates comparable to those of the centralized method, while exhibiting lower variance than Meta-SW. Notably, prior analyses using the combined data from a larger set of sites have reported a non-significant average treatment effect (Mayer et al.,, 2020; Colnet et al.,, 2024), which is consistent with our findings on this data subset.
6 Conclusion, limitations and future work
We present a theoretically grounded framework for federated causal inference from decentralized observational data. We introduce two methods for federating local propensity scores—via estimated membership probabilities or density ratios—both of which are designed to be robust to covariate shift and heterogeneity in propensity scores. Our methods improve covariate overlap, leading to more stable and accurate estimates of the ATE. The framework is flexible, allowing the integration of external control arms and supporting heterogeneous local propensity score estimation strategies.
One of the common benefits of federated learning is the ability to increase the overall sample size by leveraging data from multiple sites. While our approach does offer this advantage, it nonetheless requires a sufficiently large amount of data per site, especially when the number of variables is high. Indeed, accurate estimation of key quantities—local propensity scores, parameters of the distribution of the features per site and membership probabilities—relies on sufficient local data. For the latter, the issue becomes more pronounced as the number of sites grows.
RCTs and observational data each have strengths and limitations. Combining them can harness their respective advantages while mitigating weaknesses—supporting goals like improving heterogeneous effect estimation, addressing unobserved confounding, and enhancing generalizability. Estimating treatment effects from diverse sources is thus a promising direction for future federated causal inference. Finally, estimating the CATE in a federated setting is crucial for advancing beyond average treatment effects, towards more personalization in treatment decision-making.
Acknowledgments
This work has been done in the frame of the PEPR SN SMATCH project and has benefited from a governmental grant managed by the Agence Nationale de la Recherche under the France 2030 programme, reference ANR-22-PESN-0003. Rémi Khellaf was supported by an AXIAUM fellowship.
Appendix A Proofs
A.1 ATE Identification and Sufficient Conditions
Proof.
Under Assumptions 1 and 2, the variables in $X_{i}$ are a sufficient adjustment set, so that the ATE is identifiable as
$\tau$ $=\mathbb{E}\left[Y(1)-Y(0)\right]$
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X\right]-\mathbb{E}\left[Y(0)\mid X\right]\right]$
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,W=1\right]-\mathbb{E}\left[Y(0)\mid X,W=0\right]\right]$ Assumptions 1 and 2
$=\mathbb{E}\left[\mathbb{E}\left[Y\mid X,W=1\right]-\mathbb{E}\left[Y\mid X,W=0\right]\right]$ SUTVA
$=\mathbb{E}\left[\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\right]$ Assumption 3
A.2 Proof of Theorem 1
Proof.
Unbiasedness. We have:
$\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\cdot\mathbb{1}_{[H_{i}=k]}\mid n_{1},\dots,n_{K}\right]\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[n_{k}\right]\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$
$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$ (6)
Let us focus on the first term of the sum:
$\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]$ $=\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid H_{i}=k\right]$ SUTVA
$=\mathbb{E}\left[\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid X_{i},H_{i}=k\right]\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[W_{i}\mid X_{i},H_{i}=k\right]\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$ Assumptions 1 and 2
$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$
$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{\mathbb{E}\left[e_{k}(X_{i})\mid X_{i},H_{i}=k\right]}\right]$ Definition of $e(X_{i})$
$=\mathbb{E}\left[\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]\mathbb{E}\left[\frac{e_{k}(X_{i})}{\mathbb{E}\left[e_{k}(X_{i})\right]}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[Y_{i}(1)\mid H_{i}=k\right]$
Similarly, we have $\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]=\mathbb{E}\left[Y_{i}(0)\mid H_{i}=k\right]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
For the AIPW estimator, following the same steps as in section A.2, we have
$\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\sum_{k=1}^{K}\rho_{k}\Bigg{(}\mathbb{E}\left[\mu_{1}(X_{i})-\mu_{0}(X_{i})\mid H_{i}=k\right]$
$\quad\quad+\mathbb{E}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\mid H_{i}=k\right]\Bigg{)}$
$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\tau(X_{i})\mid H_{i}=k\right]+0+0\right)$
$=\tau$
Variance. As we consider uniformly bounded potential outcomes, that is $\forall w\in\{0,1\}$, $\exists(L,U)\in\mathbb{R}^{2},L<Y(w)<U$, and Assumption 3, we have that $\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]<\infty$ and $\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]<\infty$, so the quantities that follow are well defined:
$\mathbb{V}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]$ $=\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{V}\left[\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right]$ i.i.d. observations
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]-\tau^{2}\right)$ unbiasedness
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\right)^{2}\right]+\mathbb{E}\left[\left(\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]-2\mathbb{E}\left[0\right]-\tau^{2}\right)$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\right)^{2}\right]+\mathbb{E}\left[\left(\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]-\tau^{2}\right)$
$\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\right)^{2}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\mid X_{i}\right)^{2}\right]\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[\left(W_{i}Y_{i}\mid X_{i}\right)^{2}\right]}{e(X_{i})^{2}}\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[W_{i}\mid X_{i}\right]\mathbb{E}\left[Y_{i}^{2}\mid X_{i},W_{i}=1\right]}{e(X_{i})^{2}}\right]$ Assumption 1
$=\mathbb{E}\left[\frac{\mathbb{E}\left[Y_{i}(1)^{2}\mid X_{i}\right]}{e(X_{i})}\right]$ SUTVA
$=\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]$
Similarly, $\mathbb{E}\left[\left(\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]=\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]$. Then, the variance of the oracle multi-site IPW estimator is
$\mathbb{V}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\frac{1}{n}\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]+\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]-\tau^{2}\right).$
For the variance of the oracle multi-site centralized AIPW, we first notice that since $\mathbb{E}\left[Y_{i}(w)-\mu_{w}(X_{i})\mid X_{i}\right]=0$ for treatment $w$, we have
$A$ $=\mathrm{Cov}\left(\tau(X_{i}),\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)$
$=\mathbb{E}\left[\tau(X_{i})\left(\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\mathbb{E}\left[\tau(X_{i})\mathbb{E}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\mid X_{i}\right]\right]-\mathbb{E}\left[\tau(X_{i})\mathbb{E}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[\tau(X_{i})\frac{e(X_{i})\mathbb{E}\left[Y_{i}(1)-\mu_{1}(X_{i})\mid X_{i}\right]}{e(X_{i})}\right]-\mathbb{E}\left[\tau(X_{i})\frac{(1-e(X_{i}))\mathbb{E}\left[Y_{i}(0)-\mu_{0}(X_{i})\mid X_{i}\right]}{1-e(X_{i})}\right]$
$=0$
Then,
$\mathbb{V}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{V}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\right]+\mathbb{V}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right]+2A\right)$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X_{i}))^{2}}{e(X_{i})}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X_{i}))^{2}}{1-e(X_{i})}\right)^{2}\right]\right)$
$=\frac{1}{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X_{i}))^{2}}{e(X_{i})}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X_{i}))^{2}}{1-e(X_{i})}\right)^{2}\right]\right)$
Because these variances are well defined from what precedes, the Central Limit Theorem can be applied to $\hat{\tau}^{\mathrm{IPW^{*}}}$ and $\hat{\tau}^{\mathrm{AIPW^{*}}}$, which gives the result in Theorem 1. ∎
A.3 Proof of Theorem 2
Proof.
Unbiasedness. We first prove the unbiasedness of the local IPW estimators:
$\mathbb{E}\left[\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\mid H_{i}=k\right]$ $=\mathbb{E}\left[\frac{W_{i}Y_{i}}{e_{k}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]$ i.i.d.
$=\mathbb{E}\left[\frac{W_{i}Y_{i}{(1)}}{e_{k}(X_{i})}-\frac{(1-W_{i})Y_{i}{(0)}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]$ SUTVA
$=\mathbb{E}\left[\mathbb{E}\left[\frac{W_{i}Y_{i}{(1)}}{e_{k}(X_{i})}\mid H_{i}=k,X_{i}\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}{(0)}}{1-e_{k}(X_{i})}\mid X_{i},H_{i}=k,X_{i}\right]\right]$
$=\mathbb{E}\Bigg{[}\frac{\mathbb{E}\left[W_{i}\mid H_{i}=k,X_{i}\right]\mathbb{E}\left[Y_{i}{(1)}\mid H_{i}=k,X_{i}\right]}{e_{k}(X_{i})}$
$\quad\quad-\frac{\mathbb{E}\left[(1-W_{i})\mid H_{i}=k,X_{i}\right]\mathbb{E}\left[Y_{i}{(0)}\mid H_{i}=k,X_{i}\right]}{1-e_{k}(X_{i})}\Bigg{]}$ Ass. 1
$=\mathbb{E}\left[\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid H_{i}=k,X_{i}\right]\right]$
$=\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid H_{i}=k\right]$
$=\tau_{k}$
This yields
$\mathbb{E}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ $=\mathbb{E}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\frac{n_{k}}{n}\mathbb{E}\left[\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\mid H_{i}=k\right]\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\frac{n_{k}}{n}\right]\tau_{k}$
$=\sum_{k=1}^{K}\rho_{k}\tau_{k}$
$=\tau$
Variance. For the variance of the local ATEs, we follow the same steps as in the previous proof which yields
$\mathbb{V}\left[\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\mid H=k\right]$ $=\frac{1}{n_{k}}\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau_{k}^{2}\right)$
$=\frac{1}{n_{k}}V_{k,i}$
with $V_{k,i}=\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau_{k}^{2}$. Finally, by Lemma 2, we have
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ $=\mathbb{E}\left[\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\mid H=k\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\mid H=k\right]\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k,i}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right]$
The proof for the meta AIPW estimator follows the same steps. ∎
A.4 Decomposition of the global propensity score
By the law of total probabilities,
$e(x)$ $=\mathbb{P}(W_{i}=1\mid X_{i})$
$=\sum_{k=1}^{K}\mathbb{P}(W_{i}=1\cap H_{i}=k\mid X_{i}=x)$
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})\mathbb{P}(W_{i}=1\mid X_{i}=x,H_{i}=k)$
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})e_{k}(X_{i})$ (Eq. 3)
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\frac{P(X_{i}\mid H_{i}=k)}{P(X_{i})}e_{k}(X_{i})$ (Eq. 4)
A.5 Proof of Theorem 3
Proof.
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}$ $=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}}$
$=\sum_{k=1}^{K}\frac{n_{k}}{n}\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)$
$=\hat{\tau}^{\mathrm{IPW^{*}}}$
We prove similarly that $\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\hat{\tau}^{\mathrm{AIPW^{*}}}$. ∎
A.6 Proof of Theorem 4
We start by two technical lemmas.
Lemma 1.
Under Assumptions 1 and 2 we have,
$\mathbb{V}\left[\tau(X_{i})\right]$ $=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\right]$
Proof.
$\mathbb{V}\left[\tau(X_{i})\right]$ $=\mathbb{E}\left[\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\tau(X_{i})\mid H_{i}=k\right]\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\sum_{k=1}^{K}\mathbb{1}_{[H_{i}=k]}\tau_{H_{i}}\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\sum_{k=1}^{K}\mathbb{1}_{[H_{i}=k]}\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\right]$
Lemma 2.
For the general form of meta-analysis estimator $\hat{\tau}^{\mathrm{meta}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}$, we have
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\frac{1}{n}\mathbb{V}\left[\tau_{H_{i}}\right]$
where $V_{k}=\mathbb{V}\left[\hat{\tau}_{k}\mid H_{i}=k\right]$ is the within-site variance of the ATE estimator in site $k$ and $\mathbb{V}\left[\tau_{H_{i}}\right]=\mathbb{V}\left[\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid H_{i}\right]\right]$ is the between-sites variance of the local ATEs.
Proof.
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta}}\right]$ $=\mathbb{V}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}\right]$
$=\mathbb{E}\left[\mathbb{V}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}\mid H_{1},\dots,H_{n}\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}\mid H_{1},\dots,H_{n}\right]\right]$
$=\mathbb{E}\left[\sum_{k=1}^{K}\frac{n_{k}}{n^{2}}V_{k}\right]+\mathbb{V}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\tau_{k}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\mathbb{V}\left[\frac{1}{n}\sum_{k=1}^{K}\sum_{i=1}^{n}\mathbb{1}_{[H_{i}=k]}\tau_{H_{i}}\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\tau_{H_{i}}\sum_{k=1}^{K}\mathbb{1}_{[H_{i}=k]}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\tau_{H_{i}}\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right],$
We can now prove Theorem 4.
Proof.
We begin with IPW estimators, and then move to AIPW.
IPW.
First, applying Jensen’s inequality with the strictly convex function $t\mapsto\frac{1}{t}$ in ]0;1[]0;1[] 0 ; 1 [ and summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$, we have
$\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]<\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
if $\exists(k,k^{\prime})\in[K],e_{k}(X_{i})\neq e_{k^{\prime}}(X_{i})$, and equality if $\forall k\in[K],e_{k}(X_{i})=e(X_{i})$, i.e. if the local propensity scores are all equal to one another. Then, with the same condition on strictness and equality,
$\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]$ $\leq\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}\mid X_{i}\right]\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid X_{i}\right]\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid H_{i}=k\right]$
Similarly, $\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]\leq\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]$. Then,
$\mathbb{V}\left[\hat{\tau}^{\mathrm{fed\text{-}IPW^{*}}}\right]$ $=\frac{1}{n}\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}+\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]-\tau^{2}\right)$
$\leq\frac{1}{n}\left(\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}+\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau^{2}\right)$
$\leq\frac{1}{n}\Biggl{(}\sum_{k=1}^{K}\rho_{k}\underbrace{\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}+\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau_{k}^{2}\right)}_{:=V_{k}}+\underbrace{\sum_{k=1}^{K}\rho_{k}\tau_{k}^{2}-\tau^{2}}_{\mathbb{V}(\tau_{H})}\Biggr{)}$
On the other hand, by Lemma 2,
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right].$
Hence $\mathbb{V}\left[\hat{\tau}^{\mathrm{fed\text{-}IPW^{*}}}\right]=\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ if $\forall k\in[K],e_{k}=e$, and $\mathbb{V}\left[\hat{\tau}^{\mathrm{fed\text{-}IPW^{*}}}\right]<\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ if $\exists(k,k^{\prime})\in[K]^{2},e_{k}\neq e_{k^{\prime}}$.
AIPW.
Similarly to the IPW case, we have
$\mathbb{E}\left[\left(\frac{W_{i}(Y_{i}-\mu_{1})}{e(X_{i})}\right)^{2}\right]$ $\leq\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\left(\frac{W_{i}(Y_{i}-\mu_{1})}{e_{k}(X_{i})}\right)^{2}\mid H_{i}=k\right],$
$\mathbb{E}\left[\left(\frac{(1-W_{i})(Y_{i}-\mu_{0})}{1-e(X_{i})}\right)^{2}\right]$ $\leq\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\left(\frac{(1-W_{i})(Y_{i}-\mu_{0})}{1-e_{k}(X_{i})}\right)^{2}\mid H_{i}=k\right],$
and with Lemma 1:
$\mathbb{V}\left[\tau(X_{i})\right]=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\right].$
Then, using Lemma 2, we have the desired result. ∎
A.7 Proof of Theorem 5
Proof.
Let $f:t\mapsto\frac{1}{t(1-t)}$ with $t\in]0,1[$. $f$ is convex. Then by Jensen’s inequality with summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$,
$\mathcal{O}_{\mathrm{global}}$ $=\mathbb{E}\left[f\left(\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)\right)\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)f(e_{k}(X))\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k\mid X_{i}]}\right]f(e_{k}(X))\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}f(e_{k}(X))\mid X_{i}\right]\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}f(e_{k}(X))\right]$
$\leq\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\mathbb{E}\left[f(e_{k}(X))\mid H_{i}=k\right]$
$\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}\qed$
Appendix B Federated learning of membership weights with logistic regression
We detail how to estimate membership weights using a multinomial logistic regression model trained via Federated Averaging (FedAvg) McMahan et al., (2017), without requiring access to individual-level data.
Definition 4 (Multinomial Logistic Model).
The multinomial logistic regression model is defined as
$\omega_{k}(X_{i})=\frac{\exp(\theta_{k}^{\top}X_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\theta_{k^{\prime}}^{\top}X_{i})},$ (7)
where $\theta_{k}$ are the parameters of the model for class $k$.
We introduce the following notations: $\Theta=(\theta_{1},\cdots,\theta_{K})\in\mathbb{R}^{d\times K}$ is the matrix containing the parameters of the multinomial logistic regression, and $\ell(\Theta;\mathcal{D})$ is the negative log-likelihood of the model on a dataset $\mathcal{D}$ given by $\ell(\Theta;\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{k=1}^{K}\mathbb{1}_{\{H_{i}=k\}}\log(\omega_{k}(X_{i}))$ the softmax probabilities. $H_{i}^{\mathrm{enc}}$ is the one-hot encoding of the variable $H_{i}$.
Algorithm 1 presents the FedAvg procedure applied to training a multinomial logistic regression model for estimating membership weights.
Algorithm 1 Federated Multinomial Logistic Regression with FedAvg for Site Membership
1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ rounds of communication, $B$ batch size
2:Initialize model parameters $\Theta_{0}\in\mathbb{R}^{d\times K}$
3:for $t=1$ to $T$ do
4:     for each client $k\in[1,\dots,K]$ in parallel do
5:         $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}^{(k)}$)
6:     end for
7:     $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t}^{(k)}$ // Federated Averaging
8:end for
9:LocalUpdate($k$, $\Theta^{(k)}$, $B$):
10:for $e=1$ to $E$ do
11:     $\mathcal{B}_{k}\leftarrow$ a random batch of $B$ samples from $\mathcal{D}_{k}$
12:     $\omega_{l}(x_{i})\leftarrow\frac{\exp(\theta_{l}^{(k)}{}^{\top}x_{i})}{\sum_{l^{\prime}=1}^{K}\exp(\theta_{l^{\prime}}^{(k)}{}^{\top}x_{i})}$ // Multinomial probabilities
13:     $P_{i}\leftarrow(\omega_{1}(X_{i}),\dots,\omega_{K}(X_{i}))$
14:     $\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})\leftarrow\frac{1}{B}\sum_{i\in\mathcal{B}_{k}}x_{i}(P_{i}-H_{i}^{\mathrm{enc}})$ // Multinomial gradient
15:     $\Theta^{(k)}\leftarrow\Theta^{(k)}-\eta\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})$
16:end for
17:return $\Theta^{(k)}$
With a suitable choice of learning rate $\eta$, and a small number of local steps $E$ per round, Algorithm 1 produces parameter estimates for the multinomial logistic regression model that converge to their centralized counterparts as the number of communication rounds $T$ goes to infinity Stich, (2019); Khaled et al., (2020); Li et al., (2019). After $T$ rounds, Algorithm 1 yields an estimate $\widehat{\Theta}^{\mathrm{fed}}=(\hat{\theta}_{1}^{\mathrm{fed}},\cdots,\hat{\theta}_{K}^{\mathrm{fed}})$ of the multinomial logistic parameters which can be used to estimate the membership weights as:
$\hat{\omega}_{k}(X_{i})=\frac{\exp(\hat{\theta}_{k}^{\mathrm{fed}}{}^{\top}x_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\hat{\theta}_{k^{\prime}}^{\mathrm{fed}}{}^{\top}x_{i})}.$
Appendix C Simulation Details
[…]
Parameter Center 1 Center 2 Center 3
$d$ 10
$\mu_{1}(X)$ $\sum_{j=1}^{5}\frac{j}{10}X_{j}^{2}+\sum_{j=6}^{10}\frac{j}{10}X_{j}+X_{9}*X_{10}$
$\mu_{0}(X)$ $\sum_{j=1}^{5}\frac{3j-10}{30}X_{j}^{2}+\sum_{j=6}^{10}\frac{3j-10}{30}X_{j}+X_{1}*X_{10}$
$e_{k}$ $\mathrm{Logistic}(x,\gamma_{k})$
$\gamma_{k}$ [.$\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$ No overlap: not logistic (only controls) Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$ Good overlap: $\gamma_{2}^{\mathrm{(good)}}$ [.$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
Table 1: Common simulation parameters.
DGP A-specific settings are shown in Table 2, where $J_{d}$ is the $d\times d$ matrix of ones, and $I_{d}$ is the $d\times d$ identity matrix.
Parameter Center 1 Center 2 Center 3
$n_{k}$ 2000
$\mathcal{D}_{k}$ $\mathcal{N}(\mu_{k},\Sigma_{k})$
$\mu_{k}$ $(1,\dots,1)\in\mathbb{R}^{d}$ $(1.5,\dots,1.5)\in\mathbb{R}^{d}$ $(3,\dots,3)\in\mathbb{R}^{d}$
$\Sigma_{k}$ $I_{d}+0.5J_{d}$ $0.6I_{d}+0.4J_{d}$ $3I_{d}+0.3J_{d}$
Table 2: Simulation parameters specific to DGP A.
DGP B-specific settings are shown in Table 3.
Parameter Value
$n$ 6000
$\mathcal{D}$ $\frac{2}{3}\mathcal{N}(\mu_{1},\Sigma_{1})+\frac{1}{3}\mathcal{N}(\mu_{2},\Sigma_{2})$
$\mu_{1}$ $(0,\dots,0)\in\mathbb{R}^{d}$
$\mu_{2}$ $(1.5,\dots,1.5)\in\mathbb{R}^{d}$
$\Sigma_{1}$ $I_{d}$
$\Sigma_{2}$ $I_{d}+0.5J_{d}$
$\mathbb{P}(H_{i}=k\mid X)$ $\mathrm{Logistic}(x,\theta_{k})$
$\theta_{1}$ $[-0.5,-0.5,0.2,-0.5,-0.5,0.2,-0.5,-0.5,0.2,0.2]$
$\theta_{2}$ $[0.5,0.5,0.2,0.5,0.5,0.2,0.5,0.5,0.2,0.5]$
$\theta_{3}$ $[1,1,0.2,0.2,0.2,0.2,0.2,0.2,0.2,0.2]$
Table 3: Simulation parameters specific to DGP B.
Appendix D Overlap Improvement
Echoing theorem 5, Figures 5 and 6 display the empirical distributions (on a log-scale) of the local propensity score in site 2 ($e_{2}$) and of the global propensity score $e$ in the Poor local overlap scenario (see Section 5 for corresponding results). A good overlap is when both the propensity score distributions for the treated and control overlap, and are far from 0. We see that the poor overlap at site 2 (Figure 5), with values of $e_{2}$ on the local data close to 0, is significantly improved at the global level (Figure 6).
Figure 5: Local overlap in site 2 for the Poor local overlap scenario (DGP A).
Figure 6: Global overlap for the Poor local overlap scenario (DGP A). We see a clear improvement compared to the local overlap in site 2 (Figure 5).